# Composing Codensity Bisimulations

### v1

@@ Abstract
In addition, we propose a composition of Komorida et al.’s codensity games for bisimilarities. A novel feature of this composition is that it can also compose game invariants, which are subsets of winning positions. Under our sufficient condition of the liftability of distributive laws, composed games give an alternative proof of the preservation of bisimilarities under the composition.
@@ 1 Introduction
The results of Turi and Plotkin specifically apply to strong bisimilarity, and not directly to other notions such as similarity or behavioral metrics. To prove compositionality in these cases, the key observation is that the task is to define the composition operation at hand not only on state-based systems but also on relations (or, e.g., pseudometrics). This idea has been formalized in the theory of coalgebras by requiring a lifting of the distributive law that models the composition operator to a category of relations. The lifting of the behavior functor $F$ then specifies the notion of bisimilarity at hand, and the lifting of the syntax functor $T$ the way that relations on components can be combined into bisimulations on the composite system. It has been shown in [6] at a high level of generality that the existence of such liftings, referred to in this paper as liftability, ensures compositionality in this way. The generality there is offered by the use of liftings of the behavior functor in fibrations which goes back to [17], and which allows us to study not only relations but also, for instance, coinductively defined metrics or unary predicates (e.g., [23, 29, 16, 5]). But the main challenge here is to prove liftability; this is what we address in the current paper (Question 3.11.2).
@@
We further study a composition of codensity games. A key observation is that our composite codensity games consist of positions that are tuples of positions of codensity games for component systems. This design of the composite game directly leads to the compositionality of game invariants, which characterize winning positions. Assuming our sufficient condition of the liftability of distributive laws, we present an alternative proof of the preservation of bisimilarities along compositions (the inequality in Corollary 3.10).
@@
There is a wide range of results in the process algebra on compositionality and rule formats to prove it, which are usually focused on specific models (see, e.g., [27] for an overview). A full account is beyond the scope of this paper, which focuses instead on generality in the type of models and the type of coinductive predicates. Concerning general frameworks for compositionality in the theory of coalgebras, we have already mentioned Turi and Plotkin’s abstract GSOS format; the main innovation in the current paper is that we go beyond bisimilarity by employing the codensity lifting.
@@
Codensity lifting is first introduced to lift monads across fibrations [21]. Its instances include the Kantorivich metric on Giry monad. Later, in [28], it is extended to lift endofunctors across fibrations, and is shown to be a generalization of Baldan et al.’s Kantorovich lifting [1].
@@ 2 Preliminaries
For two natural transformations $\alpha:F\to G,\beta:G\to H$, their vertical composition is denoted by $\beta\mathbin{\bullet}\alpha$, which is the componentwise composition of $\beta$ and $\alpha$. For a functor $p:\mathbb{E}\to\mathbb{B}$ and an object $X\in\mathbb{B}$, the fiber category over $X$, denoted by $\mathbb{E}_{X}$, is the category whose objects are $P\in\mathbb{E}$ such that $pP=X$, and whose morphisms are $f$ such that $pf=id_{X}$.
@@ 2.1 ${\mathbf{CLat}_{\sqcap}}$-fibration
Examples of ${\mathbf{CLat}_{\sqcap}}$-fibrations are forgetful functors from the following categories into $\mathbf{Set}$: 1) $\mathbf{Pre}$, the category of preorders and monotone functions, 2) $\mathbf{EqRel}$, the category of equivalence relations and relation-respecting functions, 3) $\mathbf{Top}$, the category of topological spaces and continuous functions, 4) $\mathbf{PMet}$, the category of pseudometric spaces and non-expansive functions, 5) $\mathbf{ERel}$, the category of sets with endorelations, and relation-respecting functions. The forgetful functors from these categories are denoted by $U_{\mathbf{Pre}}$ etc.
A ${\mathbf{CLat}_{\sqcap}}$-fibration $p:\mathbb{E}\to\mathbb{B}$ is faithful, i.e., for any $\mathbb{E}$-object $P,Q$ and $\mathbb{B}$-morphism $f:pP\to pQ$, there is at most one $\mathbb{E}$-morphism $\dot{f}$ such that $p\dot{f}=f$. When such an $\mathbb{E}$-morphism exists, we write $f:P\mathbin{\dot{\to}}Q$. For instance, when $p$ is the forgetful functor from $\mathbf{Top}$, the notation $f:(X,O_{X})\mathbin{\dot{\to}}(Y,O_{Y})$ for a function $f:X\to Y$ is equivalent to the statement: “$f$ is continuous with respect to topologies $O_{X},O_{Y}$”. From the fiberdness, $f:P\mathbin{\dot{\to}}Q$ is equivalent to the inequality $P\sqsubseteq f^{*}Q$.
@@ 2.2 Liftings
We extensively use the concept of lifting of various categorical structures in this paper. Let $p:\mathbb{E}\to\mathbb{B}$ and $q:\mathbb{F}\to\mathbb{C}$ be functors. A lifting of a functor $F:\mathbb{B}\to\mathbb{C}$ along $p,q$ (or simply $p$ when $p=q$) is a functor ${\dot{F}}:\mathbb{E}\to\mathbb{F}$ such that $q\circ{\dot{F}}=F\circ p$. Similarly, for liftings $\dot{F},\dot{G}$ of $F,G:\mathbb{B}\to\mathbb{C}$ along $p,q$ and a natural transformation $\alpha:F\Rightarrow G$, a lifting of $\alpha$ with respect to $\dot{F},\dot{G}$ is a natural transformation $\dot{\alpha}:\dot{F}\Rightarrow\dot{G}$ such that $q\circ\dot{\alpha}=\alpha\circ p$.
To extend the concept of lifting to other categorical structures, it is convenient to set-up the arrow 2-category $\mathbf{CAT}^{\to}$ given by the following data: a 0-cell is a functor $p:\mathbb{E}\to\mathbb{B}$, a 1-cell from $p:\mathbb{E}\to\mathbb{B}$ to $q:\mathbb{F}\to\mathbb{C}$ is a pair $(F:\mathbb{B}\to\mathbb{C},\dot{F}:\mathbb{E}\to\mathbb{F})$ of functors such that $\dot{F}$ is a lifting of $F$ along $p,q$, and a 2-cell from $(F,\dot{F}):p\to q$ to $(G,\dot{G})$ is a pair of natural transformations $(\alpha,\dot{\alpha})$ such that $\dot{\alpha}$ is a lifting of $\alpha$ with respect to $\dot{F},\dot{G}$. This 2-category $\mathbf{CAT}^{\to}$ has 2-products: the terminal 0-cell is $id_{1}$, and the binary product of $p_{1},p_{2}$ is the product functor $p_{1}\times p_{2}$ (we only use finite ones in this paper). The evident forgetful 2-functor is denoted by $\mathrm{cod}:\mathbf{CAT}^{\to}\to\mathbf{CAT}$. Any categorical structure expressible within this 2-category corresponds to a pair of the categorical structure and its lifting along functors (and vice versa). For instance, a monad on $p:\mathbb{E}\to\mathbb{B}$ in $\mathbf{CAT}^{\to}$ is a pair of a monad on $\mathbb{B}$ together with its lifting along $p$ (as a monad).
For the theory of coalgebraic bisimulation in Section 3.3, we often focus on the full-sub 2-category $\mathbf{CAT}^{{\mathbf{CLat}_{\sqcap}}}$ of $\mathbf{CAT}^{\to}$ obtained by restricting 0-cells to ${\mathbf{CLat}_{\sqcap}}$-fibrations. We say that a 1-cell $(F,{\dot{F}}):p\to q$ in $\mathbf{CAT}^{\mathbf{CLat}_{\sqcap}}$ is fibered if ${\dot{F}}$ preserves Cartesian morphisms, or equivalently $(Ff)^{*}({\dot{F}}P)={\dot{F}}(f^{*}P)$ holds for any $f,P$.
The action of this 2-functor on hom-categories is denoted by the functor $\mathrm{cod}_{p,q}:\mathbf{CAT}^{\to}(p,q)\to\mathbf{CAT}(\mathrm{cod}(p),\mathrm{cod}(q))$. This is a ${\mathbf{CLat}_{\sqcap}}$-fibration, hence faithful. Thus, for a natural transformation $\alpha:F\Rightarrow G$ and liftings ${\dot{F}},{\dot{G}}$ of $F,G$ respectively, there is at most one lifting $\dot{\alpha}:{\dot{F}}\Rightarrow{\dot{G}}$ of $\alpha$. When it exists, we say that $\alpha$ is liftable with respect to ${\dot{F}},{\dot{G}}$.
@@
$\cong\{\alpha:F\Rightarrow G~{}|~{}\forall P\in\mathrm{dom}(p)~{}.~{}\alpha_{pP}:{\dot{F}}P\mathbin{\dot{\to}}{\dot{G}}P\}.$
If $\alpha_{pP}:{\dot{F}}P\mathbin{\dot{\to}}{\dot{G}}P$ holds for $P$ in a subclass $C$ of objects in $\mathbb{E}$, we say that $\alpha$ is liftable on $C$.
@@ 3 Overview
The theory of coalgebras provides a categorical framework for expressing various transition systems in a unified manner. Let $\mathbb{C}$ be a category and $F:\mathbb{C}\to\mathbb{C}$ be an endofunctor (called behavior functor). An $F$-coalgebra is a pair of an object $X$ (called carrier) and a morphism $c:X\to F(X)$. A morphism from $c:X\to F(X)$ to $d:Y\to F(Y)$ is a morphism $h:X\to Y$ such that $Fh\circ c=d\circ h$.
@@
We write ${\mathcal{P}}$ for the covariant powerset functor on $\mathbf{Set}$. A ${\mathcal{P}}$-coalgebra $c:X\to{\mathcal{P}}(X)$ bijectively corresponds to a binary relation on $X$, that is, a Kripke frame.
@@
Fix an alphabet $\Sigma$. We define the endofunctor ${F_{\mathrm{da}}}$ on $\mathbf{Set}$ by ${F_{\mathrm{da}}}\triangleq 2\times(-)^{\Sigma}$. An ${F_{\mathrm{da}}}$-coalgebra $c:X\to{F_{\mathrm{da}}}(X)$ bijectively corresponds to a deterministic automata over alphabet $\Sigma$.
The category of $F$-coalgebras and morphisms between them is denoted by $\mathbf{Coalg}(F)$, and the evident forgetful functor to $\mathbb{C}$ is denoted by $U_{F}$. We note that $(\mathbf{Coalg}(F))^{N}\cong\mathbf{Coalg}(F^{N})$. When $(F,{\dot{F}}):p\to p$ is a 1-cell in $\mathbf{CAT}^{\to}$, the application of $p$ to ${\dot{F}}$-coalgebras extends to a functor $\mathbf{Coalg}(p):\mathbf{Coalg}({\dot{F}})\to\mathbf{Coalg}(F)$.
@@ 3.1 Coalgebraic Bisimulation
One of the important concepts in state transition systems is the identification of two states that behave in the same way. A classical notion of behavioral equivalence is bisimilarity, which has been extensively studied in process algebra, coalgebra theory and modal logic. Recently, the concept of bisimulation, initially formulated as a binary relation, is expressed using other spatial structures, such as pseudometrics [2] and topologies [28]. A uniform treatment of these spatial representations of bisimulation have been developed using the framework of ${\mathbf{CLat}_{\sqcap}}$-fibrations [28]. We employ the formulation of bisimulation as coalgebras in suitable fibrations. This formulation goes back to the seminal work by Hermida and Jacobs [17], and the key ingredient is the lifting of behavior functors along fibrations. We recall their theory here.
Let $p:\mathbb{E}\to\mathbb{B}$ be a ${\mathbf{CLat}_{\sqcap}}$-fibration and $c:X\to F(X)$ be an $F$-coalgebra. Hermida and Jacobs’ theory first chooses a lifting ${\dot{F}}$ of $F$ along $p$; the main body of the definition of bisimulation is pack into it. Then we formulate a ${\dot{F}}$-bisimulation on a coalgebra $c:X\to F(X)$ to be a ${\dot{F}}$-coalgebra $\dot{c}:P\to{\dot{F}}(P)$ such that $p(\dot{c})=c$ (hence $p(P)=X$). The morphism $\dot{c}$ witnesses that $P$ is respected by the underlying transition system $p(c)$. Since there exists at most one such arrow $\dot{c}$ for a given $P$, a ${\dot{F}}$-bisimulation bijectively corresponds to $P\in\mathbb{E}_{X}$ satisfying $P\sqsubseteq c^{*}\circ{\dot{F}}(P)$. We thus define the set of bisimulations on $c$ by
$\mathbf{Bisim}({\dot{F}},c)$$\triangleq\{P\in\mathbb{E}_{X}~{}|~{}P\sqsubseteq c^{*}\circ{\dot{F}}(P)\},$
and impose a partial order on it by restricting the one on $\mathbb{E}_{X}$. Since $\mathbb{E}_{X}$ is a complete lattice, it contains the greatest postfixpoint $\nu(c^{*}\circ{\dot{F}})$ corresponding to the bisimilarity.
@@
$\dot{\mathcal{P}}(X,P)\triangleq({\mathcal{P}}(X),\{(U,V)~{}|~{}$ $\forall x\in U~{}.~{}\exists y\in V~{}.~{}(x,y)\in P\wedge$
@@
Then $P$ is a $\dot{\mathcal{P}}$-bisimulation on $c:X\to{\mathcal{P}}(X)$ if and only if it satisfies the standard bisimulation condition:
@@
$\forall x^{\prime}\in c(y)~{}.~{}\exists y^{\prime}\in c(x)~{}.~{}(x^{\prime},y^{\prime})\in P.$
@@
${\dot{F}_{\mathrm{da}}}(P)\triangleq\{((t_{1},\rho_{1}),(t_{2},\rho_{2}))\mid t_{1}=t_{2}\wedge\forall a\in\Sigma~{}.~{}(\rho_{1}(a),\rho_{2}(a))\in P\}.$
Let $P$ be a ${\dot{F}_{\mathrm{da}}}$-bisimulation on $c:X\to{F_{\mathrm{da}}}(X)$. Then any pair of states $(x,y)\in P$ are language equivalent, that is, the set of words accepted from $x$ coincides with that of $y$. Moreover, ${\dot{F}_{\mathrm{da}}}$-bisimilarity coincides with the set of all language-equivalent state pairs.

### v2

@@ Abstract
In addition, we propose a compositional variant of Komorida et al.’s codensity games for bisimilarities. A novel feature of this variant is that it can also compose game invariants, which are subsets of winning positions. Under our sufficient condition of the liftability of distributive laws, composed games give an alternative proof of the preservation of bisimilarities under the composition.
@@ 1 Introduction
The results of Turi and Plotkin specifically apply to strong bisimilarity, and not directly to other notions such as similarity or behavioral metrics. To prove compositionality in these cases, the key observation is that the task is to define the composition operation at hand not only on state-based systems but also on relations (or, e.g., pseudometrics). This idea has been formalized in the theory of coalgebras by requiring a lifting of the distributive law that models the composition operator to a category of relations. The lifting of the behavior functor $F$ then specifies the notion of bisimilarity at hand, and the lifting of the syntax functor $T$ specifies the way that relations on components can be combined into bisimulations on the composite system. It has been shown in [6] at a high level of generality that the existence of such liftings, referred to in this paper as liftability, ensures compositionality. The generality there is offered by the use of liftings of the behavior functor in fibrations which goes back to [17], and which allows us to study not only relations but also, for instance, coinductively defined metrics or unary predicates (e.g., [23, 29, 16, 5]). But the main challenge here is to prove liftability; this is what we address in the current paper (Question 3.11.2).
@@
We further study a compositional variant of codensity games. A key observation is that our composite codensity games consist of positions that are tuples of positions of codensity games for component systems. This design of the composite game directly leads to the compositionality of game invariants, which characterize winning positions. Assuming our sufficient condition for the liftability of distributive laws, we present an alternative proof of the preservation of bisimilarities along compositions (the inequality in Corollary 3.10).
@@
There is a wide range of results in the process algebra literature on proving compositionality and on rule formats that guarantee it, usually focused on transition systems (see, e.g., [27] for an overview). A full account is beyond the scope of this paper. We focus instead on generality in the type of models and the type of coinductive predicate. Concerning general coalgebraic frameworks for compositionality, we have already mentioned Turi and Plotkin’s abstract GSOS format; the main innovation in the current paper is that we go beyond bisimilarity by employing the codensity lifting.
@@
Codensity liftings are first introduced to lift monads across fibrations [21]. They subsume the Kantorivich metric on the Giry monad. In [28], they are extended to lift endofunctors across ${\mathbf{CLat}_{\sqcap}}$-fibrations over $\mathbf{Set}$, and are shown to be a generalization of Baldan et al.’s Kantorovich lifting [1]. Prior to these papers, the monadic property of the Kantorovich metric is studied in an unpublished manuscript by van Breugel [34]. Codensity liftings for endofunctors can be done in other base categories (such as $\mathbf{Meas}$ and $\mathbf{Vect}$) without modification; see [23]. In [13], the Kantorovich liftings are studied in the context of Lawvere metric space, and are related with other notions of liftings, such as lax extensions and (monotone) predicate liftings.
@@ 2 Preliminaries
For two natural transformations $\alpha\colon F\to G,\beta\colon G\to H$, their vertical composition is denoted by $\beta\mathbin{\bullet}\alpha$, which is the componentwise composition of $\beta$ and $\alpha$. For a functor $p\colon\mathbb{E}\to\mathbb{B}$ and an object $X\in\mathbb{B}$, the fiber category over $X$, denoted by $\mathbb{E}_{X}$, is the category whose objects are $P\in\mathbb{E}$ such that $pP=X$, and whose morphisms are $f$ such that $pf=id_{X}$.
@@ 2.1 ${\mathbf{CLat}_{\sqcap}}$-fibration
Examples of ${\mathbf{CLat}_{\sqcap}}$-fibrations are forgetful functors from the following categories into $\mathbf{Set}$: 1) $\mathbf{Pre}$, the category of preorders and monotone functions, 2) $\mathbf{EqRel}$, the category of equivalence relations and relation-respecting functions, 3) $\mathbf{Top}$, the category of topological spaces and continuous functions, 4) $\mathbf{PMet}$, the category of pseudometric spaces and non-expansive functions, 5) $\mathbf{ERel}$, the category of sets with endorelations and relation-respecting functions. The forgetful functors are denoted by $U_{\mathbf{Pre}}$ etc.
Every ${\mathbf{CLat}_{\sqcap}}$-fibration $p\colon\mathbb{E}\to\mathbb{B}$ is faithful, i.e., for any $\mathbb{E}$-object $P,Q$ and $\mathbb{B}$-morphism $f\colon pP\to pQ$, there is at most one $\mathbb{E}$-morphism $\dot{f}$ such that $p\dot{f}=f$. When such an $\mathbb{E}$-morphism exists, we write $f\colon P\mathbin{\dot{\to}}Q$. For instance, when $p$ is the forgetful functor from $\mathbf{Top}$, the notation $f\colon(X,O_{X})\mathbin{\dot{\to}}(Y,O_{Y})$ for a function $f\colon X\to Y$ is equivalent to the statement: “$f$ is continuous with respect to topologies $O_{X},O_{Y}$”. From the fiberedness, $f\colon P\mathbin{\dot{\to}}Q$ is equivalent to the inequality $P\sqsubseteq f^{*}Q$.
@@ 2.2 Liftings
We extensively use the concept of lifting of various categorical structures in this paper. Let $p\colon\mathbb{E}\to\mathbb{B}$ and $q\colon\mathbb{F}\to\mathbb{C}$ be functors. A lifting of a functor $F\colon\mathbb{B}\to\mathbb{C}$ along $p,q$ (or simply $p$ when $p=q$) is a functor ${\dot{F}}\colon\mathbb{E}\to\mathbb{F}$ such that $q\circ{\dot{F}}=F\circ p$. Similarly, for liftings $\dot{F},\dot{G}$ of $F,G\colon\mathbb{B}\to\mathbb{C}$ along $p,q$ and a natural transformation $\alpha\colon F\Rightarrow G$, a lifting of $\alpha$ with respect to $\dot{F},\dot{G}$ is a natural transformation $\dot{\alpha}\colon\dot{F}\Rightarrow\dot{G}$ such that $q\circ\dot{\alpha}=\alpha\circ p$.
To extend the concept of lifting to other categorical structures, it is convenient to set-up the arrow 2-category $\mathbf{CAT}^{\to}$ given by the following data: a 0-cell is a functor $p\colon\mathbb{E}\to\mathbb{B}$, a 1-cell from $p\colon\mathbb{E}\to\mathbb{B}$ to $q\colon\mathbb{F}\to\mathbb{C}$ is a pair $(F\colon\mathbb{B}\to\mathbb{C},\dot{F}\colon\mathbb{E}\to\mathbb{F})$ of functors such that $\dot{F}$ is a lifting of $F$ along $p,q$, and a 2-cell from $(F,\dot{F})\colon p\to q$ to $(G,\dot{G})$ is a pair of natural transformations $(\alpha,\dot{\alpha})$ such that $\dot{\alpha}$ is a lifting of $\alpha$ with respect to $\dot{F},\dot{G}$. This 2-category $\mathbf{CAT}^{\to}$ has 2-products: the terminal 0-cell is $id_{1}$, and the binary product of $p_{1},p_{2}$ is the product functor $p_{1}\times p_{2}$ (we only use finite ones in this paper). The evident forgetful 2-functor is denoted by $\mathrm{cod}\colon\mathbf{CAT}^{\to}\to\mathbf{CAT}$. Any categorical structure expressible within this 2-category corresponds to a pair of the categorical structure and its lifting along functors (and vice versa). For instance, a monad on $p\colon\mathbb{E}\to\mathbb{B}$ in $\mathbf{CAT}^{\to}$ is a pair of a monad on $\mathbb{B}$ together with its lifting along $p$ (as a monad).
For the theory of coalgebraic bisimulation in Section 3.3, we often focus on the full-sub 2-category $\mathbf{CAT}^{{\mathbf{CLat}_{\sqcap}}}$ of $\mathbf{CAT}^{\to}$ obtained by restricting 0-cells to ${\mathbf{CLat}_{\sqcap}}$-fibrations. We say that a 1-cell $(F,{\dot{F}})\colon p\to q$ in $\mathbf{CAT}^{\mathbf{CLat}_{\sqcap}}$ is fibered if ${\dot{F}}$ preserves Cartesian morphisms, or equivalently $(Ff)^{*}({\dot{F}}P)={\dot{F}}(f^{*}P)$ holds for any $f,P$.
The action of this 2-functor on hom-categories is denoted by the functor $\mathrm{cod}_{p,q}\colon\mathbf{CAT}^{\to}(p,q)\to\mathbf{CAT}(\mathrm{cod}(p),\mathrm{cod}(q))$. This is a ${\mathbf{CLat}_{\sqcap}}$-fibration, hence faithful. Thus, for a natural transformation $\alpha\colon F\Rightarrow G$ and liftings ${\dot{F}},{\dot{G}}$ of $F,G$ respectively, there is at most one lifting $\dot{\alpha}\colon{\dot{F}}\Rightarrow{\dot{G}}$ of $\alpha$. When it exists, we say that $\alpha$ is liftable with respect to ${\dot{F}},{\dot{G}}$.
@@
$\cong\{\alpha\colon F\Rightarrow G~{}|~{}\forall P\in\mathrm{dom}(p)~{}.~{}\alpha_{pP}\colon{\dot{F}}P\mathbin{\dot{\to}}{\dot{G}}P\}.$
If $\alpha_{pP}\colon{\dot{F}}P\mathbin{\dot{\to}}{\dot{G}}P$ holds for $P$ in a subclass $C$ of objects in $\mathbb{E}$, we say that $\alpha$ is liftable on $C$.
@@ 3 Overview
The theory of coalgebras provides a categorical framework for expressing various transition systems in a unified manner. Let $\mathbb{C}$ be a category and $F\colon\mathbb{C}\to\mathbb{C}$ be an endofunctor (called behavior functor). An $F$-coalgebra is a pair of an object $X$ (called carrier) and a morphism $c\colon X\to F(X)$. A morphism from $c\colon X\to F(X)$ to $d\colon Y\to F(Y)$ is a morphism $h\colon X\to Y$ such that $Fh\circ c=d\circ h$.
@@
We write ${\mathcal{P}}$ for the covariant powerset functor on $\mathbf{Set}$. A ${\mathcal{P}}$-coalgebra $c\colon X\to{\mathcal{P}}(X)$ bijectively corresponds to a binary relation on $X$, that is, a Kripke frame.
@@
Fix an alphabet $\Sigma$. We define the endofunctor ${F_{\mathrm{da}}}$ on $\mathbf{Set}$ by ${F_{\mathrm{da}}}\coloneq 2\times(-)^{\Sigma}$. An ${F_{\mathrm{da}}}$-coalgebra $c\colon X\to{F_{\mathrm{da}}}(X)$ bijectively corresponds to a deterministic automata over alphabet $\Sigma$.
The category of $F$-coalgebras and morphisms between them is denoted by $\mathbf{Coalg}(F)$, and the evident forgetful functor to $\mathbb{C}$ is denoted by $U_{F}$. We note that $(\mathbf{Coalg}(F))^{N}\cong\mathbf{Coalg}(F^{N})$. When $(F,{\dot{F}})\colon p\to p$ is a 1-cell in $\mathbf{CAT}^{\to}$, the application of $p$ to ${\dot{F}}$-coalgebras extends to a functor $\mathbf{Coalg}(p)\colon\mathbf{Coalg}({\dot{F}})\to\mathbf{Coalg}(F)$.
@@ 3.1 Coalgebraic Bisimulation
One of the important concepts in state transition systems is the identification of two states that behave in the same way. A classical notion of behavioral equivalence is bisimilarity, which has been extensively studied in process algebra, coalgebra theory and modal logic. Recently, the concept of coalgebraic bisimulation, initially formulated as a binary relation, is expressed using other spatial structures, such as pseudometrics [2] and topologies [28]. A uniform treatment of these spatial representations of bisimulation have been developed using the framework of ${\mathbf{CLat}_{\sqcap}}$-fibrations [28]. We employ the formulation of bisimulation as coalgebras in suitable fibrations. This formulation goes back to the seminal work by Hermida and Jacobs [17], and the key ingredient is the lifting of behavior functors along fibrations. We recall their theory here.
Let $p\colon\mathbb{E}\to\mathbb{B}$ be a ${\mathbf{CLat}_{\sqcap}}$-fibration and $c\colon X\to F(X)$ be an $F$-coalgebra. Hermida and Jacobs’ theory first chooses a lifting ${\dot{F}}$ of $F$ along $p$; the main body of the definition of bisimulation is packed into it. Then we formulate an ${\dot{F}}$-bisimulation on a coalgebra $c\colon X\to F(X)$ to be a ${\dot{F}}$-coalgebra $\dot{c}\colon P\to{\dot{F}}(P)$ such that $p(\dot{c})=c$ (hence $p(P)=X$). The morphism $\dot{c}$ witnesses that $P$ is respected by the underlying transition system $p(c)$. An ${\dot{F}}$-bisimulation bijectively corresponds to $P\in\mathbb{E}_{X}$ satisfying $P\sqsubseteq c^{*}\circ{\dot{F}}(P)$. We thus define the set of bisimulations on $c$ by
$\mathbf{Bisim}({\dot{F}},c)$ :-$\coloneq\{P\in\mathbb{E}_{X}~{}|~{}P\sqsubseteq c^{*}\circ{\dot{F}}(P)\},:$
and impose a partial order on it by restricting the one on $\mathbb{E}_{X}$. Since $\mathbb{E}_{X}$ is a complete lattice, it contains the greatest postfixpoint $\nu(c^{*}\circ{\dot{F}})$ corresponding to bisimilarity.
@@
$\dot{\mathcal{P}}(X,P)\coloneq({\mathcal{P}}(X),\{(U,V)~{}|~{}$ $\forall x\in U~{}.~{}\exists y\in V~{}.~{}(x,y)\in P\wedge$
@@
Then $P$ is a $\dot{\mathcal{P}}$-bisimulation on $c\colon X\to{\mathcal{P}}(X)$ if and only if it satisfies the standard bisimulation condition:
@@
$\forall y^{\prime}\in c(y)~{}.~{}\exists x^{\prime}\in c(x)~{}.~{}(x^{\prime},y^{\prime})\in P.$
@@
${\dot{F}_{\mathrm{da}}}(P)\coloneq\{((t_{1},\rho_{1}),(t_{2},\rho_{2}))\mid t_{1}=t_{2}\wedge\forall a\in\Sigma~{}.~{}(\rho_{1}(a),\rho_{2}(a))\in P\}.$
Let $P$ be an ${\dot{F}_{\mathrm{da}}}$-bisimulation on $c\colon X\to{F_{\mathrm{da}}}(X)$. Then any pair of states $(x,y)\in P$ are language equivalent, that is, the set of words accepted from $x$ coincides with that of $y$. Moreover, ${\dot{F}_{\mathrm{da}}}$-bisimilarity coincides with the set of all language-equivalent state pairs.